OpenStudy (roadjester):

Find an equation of the tangent line to the curve x*e^y + y*e^x = 1

OpenStudy (roadjester):

At the point (0,1). Forgot about that part of the question.

OpenStudy (anonymous):

what i wrote was wrong, it should be \[e^y+xe^yy'+y'e^x+ye^x=0\]

OpenStudy (anonymous):

replace x by 0, y by 1 and solve for y'

OpenStudy (roadjester):

Just so you know, my script stopped for some reason and I have no idea what that means so I can't view anything from the equation editor or drawings. I just get "Error".

OpenStudy (anonymous):

\[e+y'+1=0\] \[y'=-e-1\] if my algebra is correct

OpenStudy (roadjester):

I've tried reload but nothing.

OpenStudy (anonymous):

refresh doesn't work?

OpenStudy (roadjester):

"Math Processing Error" Nope, it won't work.

OpenStudy (roadjester):

Shouldn't I solve for y' first, then substitute?

OpenStudy (roadjester):

I know I need to use the product rule, but how do I solve for y if y is a part of e^y (an exponential)

OpenStudy (anonymous):

the derivative of e^y with resepect to x is e^y y' by the chain rule

OpenStudy (roadjester):

Just out of curiosity, how did you get that into a pdf?